Determine which of the following functions are of bounded variation on[0,1]
a. f(x)= x^2sin(1/X) if x#0 ,f(0)=0
b. f(x)= √x sinx,if x#0,f(0)=0
1
Expert's answer
2020-05-11T08:51:23-0400

"f(x)=x^2sin(\frac1x)"

"f(0)=0;f(1)=sin(1)"

In the interval (0,1):

"0\leq sin(\frac1x)\leq 1"

"0\leq x^2sin(\frac1x)\leq x^2"

So,the function doesn't blow up to infinity,thus, f(x) is bounded.

(b)"f(x)=\sqrt{x}sin x"

"f(0)=0;f(1)=sin(1)"

"0\leq sinx \leq1"

"0\leq \sqrt{x}sinx \leq\sqrt x"

So,the function doesn't blow up to infinity,thus, f(x) is bounded.